# Variational Bayesian latent estimation with compressive autoencoders

### v2

@@ 4.3 BSD experiments
We now compare VBLE using CAEs to state-of-the-art PnP baselines on BSD dataset, which is composed of various natural images. We consider three state-of-the-art methods: DPIR [10], PnP-ADMM [8] and DiffPIR [18]. Recall that, unlike PnP methods, VBLE enables posterior sampling. Hence, we also compare to the MCMC approach PnP-ULA [33], for which $10^{5}$ MCMC iterations are performed. For DPIR, PnP-ADMM and PnP-ULA, we use an available version of DRUNet [10] trained on BSD. For DiffPIR, a pretrained model on ImageNet [62] is used, that we finetune on BSD dataset. Similarly to FFHQ results, we provide restoration results with VBLE and its deterministic counterpart using CAE (denoted by MAPz wCAE in the experiments).
BSD Deblur (Gaussian) SISR $\times$4 SISR $\times$2
Method PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$
VBLE wCAE 29.77 0.2002 0.8617 25.47 0.3459 0.6977 29.65 0.1830 0.8749
MAPz wCAE 29.28 0.2215 0.8405 25.38 0.3462 0.6966 29.65 0.1675 0.8757
PnP-ULA 27.71 0.2675 0.7809 24.82 0.3730 0.6894 28.42 0.1659 0.8763
DPIR 29.72 0.2127 0.8623 25.18 0.3652 0.6937 29.55 0.1706 0.8784
PnP-ADMM 28.96 0.2155 0.8456 24.77 0.4227 0.6659 29.31 0.1907 0.8678
DiffPIR 29.26 0.1832 0.8420 25.03 0.3164 0.6771 29.20 0.1645 0.8539
Table 2: BSD results on diverse inverse problems: Gaussian deblurring ($\sigma_{blur}=1$, $\sigma=7.65/255$), SISR ×$\times 4$, SISR ×$\times 2$. Methods in italics allow for posterior sampling.
Figure 5: Qualitative image restoration results on BSD. From top to bottom: Motion deblurring ($\sigma=7.65/255$), SISR ×$\times 4$, SISR ×$\times 2$.
Tab. 2 shows quantitative results on deblurring and SISR inverse problems, while Fig. 5 provides a visual comparison. VBLE exhibits very consistent performance as it reaches the first or second rank for a majority of inverse problems and metrics. Besides, VBLE clearly outperforms PnP-ULA, which is the only other posterior sampling method. Visually, except for PnP-ULA, all methods perform very well. In particular, VBLE shows very relevant results, staying particularly faithful to the image. Additional inpainting with structured masks results are given in Fig. 6 together with computation times. The latter shows that VBLE wCAE is as fast as DiffPIR, slightly slower than DPIR, and several orders of magnitude faster than the other posterior sampling baseline PnP-ULA.
DPIR 3s
DiffPIR 21s
PnP-
ULA
1h23m29s
VBLE
24s
MAPz
wCAE
23s
Figure 6: Left: Inpainting results on BSD, with Gaussian noise $\sigma=2.55/255$. Right: Computational load for the different methods. Computation time is given for inpainting one BSD image on a Nvidia Quadro RTX 8000 GPU.
@@ 4.4 Posterior distribution quality assessment
Sections 4.3 and 4.2 show that VBLE using CAEs reaches state-of-the-art point estimation results for image restoration, while outperforming the MCMC method PnP-ULA. Another keypoint of VBLE is its ability to derive posterior samples and uncertainties very efficiently as variational inference enables to explicitly approximate the true posterior. Thus, we analyze in this section the quality of VBLE approximate posterior distribution.
Figure 7: Visual comparison between VBLE and PnP-ULA on several inverse problems. From top to bottom: Motion deblurring ($\sigma=7.65/255$), SISR ×$\times 2$, block inpainting ($\sigma=2.55/255$).
Figure 7 shows a visual comparison of VBLE and PnP-ULA posterior samples and marginal standard deviations, estimated from posterior sampling with 100 samples. Marginal standard deviations of VBLE seem relevant and their maximum value is generally close to PnP-ULA deviations. Then for low frequency zones, VBLE exhibits low deviations in contrast to PnP-ULA. Therefore, VBLE also exhibits realistic posterior samples, while PnP-ULA ones exhibit more noise. VBLE posterior sampling ability is shown in more detail in Fig. 8. Posterior samples are sharp and realistic, showing the interest of formulating the posterior distribution in the latent space, with a uniform distribution. The MMSE-x estimate is smoother than the samples and, hence, tends to be more faithful to the target image.
@@
Figure 8: Example of VBLE posterior sampling ability. The MMSE-x, which is the estimate used in all quantitative comparison, corresponds to the average over samples in the image space.
@@ 5 Conclusion
In this paper, we have proposed the use of variational compressive autoencoders for regularizing imaging inverse problems. Additionally, we have introduced the VBLE algorithm which performs variational inference in the latent space of CAEs or VAEs. VBLE enables efficient approximation of the posterior distribution associated with image restoration tasks. In our experiments, compressive autoencoders, combined with VBLE, achieve state-of-the-art results and provide compelling posterior sampling abilities. These demonstrate the interest of CAEs for image restoration as well as the possibility to estimate simply and efficiently the posterior distribution in the latent space of a generative model.

### v4

@@ 4.3 BSD experiments
Figure 5: Qualitative image restoration results on BSD. From top to bottom: Motion deblurring ($\sigma=7.65/255$), SISR $\times 4$, SISR $\times 2$ and two block inpainting experiments ($\sigma=2.55/255$).
1 solution 100 post.samples GPU mem. #NN #Backward
Time #NFEs Time #NFEs (MiB) params
VBLE wCAE 22.4s 1000 24.3s 1 100 400 29M 1000
PnP-ULA x x 1h23min 100 000 341 32M 0
DPIR 1.9s 50 x x 280 32M 0
PnP-ADMM 4.1s 150 x x 282 32M 0
DiffPIR 21s 100 35min 10 000 2690 552M 0
DDRM 44.5s 100 1h14min 10 000 36 270 552M 0
Table 3: Complexity analysis of our method and the baselines of Section IV.C-D. # means ”number of”. The times, the number of Neural Function Evaluations (NFEs), the number of backward steps, and maximum GPU memory are given for inpainting a BSD image on a Nvidia Quadro RTX 8000. Time and #NFEs are given for computing either 1 estimate or 100 posterior samples (for VBLE, optimization AND sampling included).
We now compare VBLE using CAEs to state-of-the-art PnP baselines on BSD dataset, which is composed of various natural images. We consider four state-of-the-art methods: DPIR [70], PnP-ADMM [66], DiffPIR [73] and DDRM [42]. As for FFHQ, we do not perform motion blur with DDRM as it is not implemented in the source code. We also compare to the MCMC approach PnP-ULA [44], which is a state-of-the-art posterior sampling method. For DPIR, PnP-ADMM, and PnP-ULA, we use an available version of DRUNet [70] trained on BSD. For DiffPIR and DDRM, a pretrained model on ImageNet [24] is used, that we finetune on BSD dataset. For PnP-ULA, $10^{5}$ MCMC iterations are performed. Similarly to FFHQ results, we provide restoration results with VBLE and its deterministic counterpart using CAE (denoted by MAPz wCAE in the experiments).
Table 2 shows quantitative results on deblurring and SISR inverse problems, while Fig. 5 provides a visual comparison. VBLE exhibits very consistent performance as it reaches the first or second rank for a majority of inverse problems and metrics. Besides, VBLE clearly outperforms PnP-ULA. Visually, except for PnP-ULA, all methods perform very well. In particular, the diffusion-based results are particularly sharp, while VBLE shows very relevant results, staying faithful to the image.
Additionally, Tab. 3 provides insights about the computation time, maximum GPU memory, and other complexity-related information. VBLE time includes the variational inference optimization and the sampling for one restoration. In particular, for drawing 100 posterior samples, VBLE is much faster than DiffPIR, DDRM, and PnP-ULA, with a lower computational load than DiffPIR and DDRM. Note that algorithm 1 is performed during 1000 iterations, thus VBLE, even for computing a single estimate, requires 1000 NFEs.
Therefore, VBLE yields state-of-the-art image restoration results compared to very competitive diffusion and PnP baselines while outperforming the posterior sampling methods in terms of computation time and GPU load. This demonstrates the huge interest of VBLE with CAE for image restoration tasks, especially when uncertainty quantification is required as it enables posterior sampling with a far lower computational cost than MCMC or diffusion-based methods.
@@ 4.4 Posterior distribution quality assessment
FFHQ Deblur (Gaus.) $\sigma_{k}=1$ Deblur (Motion) SISR $\times 2$ SISR $\times 4$
Method Time PSNR $\uparrow$ LPIPS $\downarrow$ ICP 95% PSNR $\uparrow$ LPIPS $\downarrow$ ICP 95% PSNR $\uparrow$ LPIPS $\downarrow$ ICP 95% PSNR $\uparrow$ LPIPS $\downarrow$ ICP 95%
VBLE wCAE 27sec 34.48 0.1473 0.85 32.22 0.1933 0.87 36.16 0.1251 0.57 31.18 0.2290 0.73
DiffPIR-100 8min10sec 34.78 0.1445 0.76 32.95 0.1869 0.75 35.78 0.1207 0.79 31.21 0.2252 0.69
PnP-ULA 1h23min 32.51 0.2014 0.99 30.35 0.2367 0.99 30.82 0.1570 0.92 28.94 0.2495 0.68
BSD Deblur (Gaus.), $\sigma_{k}=1$ Deblur (Motion) SISR $\times 2$ SISR $\times 4$
Method Time PSNR $\uparrow$ LPIPS $\downarrow$ ICP 95% PSNR $\uparrow$ LPIPS $\downarrow$ ICP 95% PSNR $\uparrow$ LPIPS $\downarrow$ ICP 95% PSNR $\uparrow$ LPIPS $\downarrow$ ICP 95%
VBLE wCAE 27sec 29.62 0.1901 0.81 28.38 0.2356 0.78 29.55 0.1802 0.44 25.53 0.3507 0.60
DiffPIR-100 35min00sec 29.88 0.1847 0.67 29.09 0.2351 0.70 29.61 0.1789 0.61 25.36 0.3394 0.69
PnP-ULA 1h23min 27.71 0.2675 0.95 27.17 0.2643 0.98 28.42 0.1659 0.78 24.82 0.3730 0.59
Table 4: Comparison of posterior sampling methods. DiffPIR-100 corresponds DiffPIR, which is run 100 times for each image to get 100 posterior samples. The time is the computation time needed for drawing 100 posterior samples for an inpainting problem on a Nvidia Quadro RTX 8000 (for VBLE time, optimization AND sampling is included). ICP 95% is the confidence coverage probability of level 95% (the closer to 0.95, the better). $\sigma=7.65$ for deblurring.
Figure 6: Visual comparison of posterior sampling ability of VBLE, DiffPIR and PnP-ULA. Top: SISR $\times 2$ on FFHQ. Bottom: Motion deblurring on BSD. The MMSE estimate is the average of 100 samples for each method. The true error denotes the absolute error |xx^||x-$\hat${x}| between the ground truth and the MMSE. The predicted error (95%) denotes the 95% quantile of $\{|\hat{x}-x_{i}|\}_{i}$ with $x_{i}$ a posterior sample.
@@
Figure 7: Example of VBLE posterior sampling ability. Left: The MMSE-x, which is the estimate used in all quantitative comparison, corresponds to the average over samples in the image space. Marginal std corresponds to pixelwise marginal standard deviations computed with 100 posterior samples. Right: Illustration of the latent diversity on FFHQ for a SISR $\times 4$ problem. The samples lie on a grid. The coordinates correspond to the latent coordinates of the samples, projected on two orthogonal directions with high variance in the latent space.
Sections 4.2 and 4.3 show that VBLE using CAEs reaches state-of-the-art point estimation results for image restoration while outperforming the other posterior sampling methods. In this section, we analyze the quality of VBLE approximate posterior distribution, qualitatively and quantitatively compared to DiffPIR and PnP-ULA. First, Tab. 4 provides an in-depth comparison between VBLE, DiffPIR, and PnP-ULA. In this table, the metrics are computed from the posterior mean, estimated by the average of 100 posterior samples for each image. Thus, DiffPIR is run 100 times for each of these images. Finally, the different metrics are averaged on the whole test set. DiffPIR has very good PSNR and LPIPS but is far slower than VBLE. Furthermore, in this table, we evaluate the quality of the different posteriors using confidence intervals. Precisely, for each pixel of an image, we compute a confidence interval of a given level using 100 posterior samples. Then, we evaluate the interval coverage probability (ICP), that is the proportion of ground truth pixels lying in the predicted interval. In Tab. 4, ICPs at level $95\%$ are provided for several inverse problems. PnP-ULA has the best ICPs, while VBLE ICPs are slightly below and DiffPIR is outperformed by PnP-ULA and VBLE. Hence, VBLE offers an interesting trade-off as it remains competitive with DiffPIR, while having a relevant approximate posterior and above all, being significantly faster than both methods for posterior sampling.
Second, in Fig. 6, we provide, for the different methods, some posterior samples, as well as a comparison between the true pixelwise error and the 95% quantile of the predicted error. A coverage graph is also provided for each image. Precisely, for each image, a confidence interval of level $\beta$ is derived from the squared error for each pixel. Then, the empirical coverage is plotted as a function of the theoretical coverage $\beta$ varying from 0 to 100%. This evaluation process is the pixelwise counterpart of the one proposed in [54]. For both images of Fig. 6, VBLE exhibits on average better coverage graphs than DiffPIR and PnP-ULA. Visually, the predicted errors are relevant. In comparison, PnP-ULA predictions are over or underestimated depending on the problems, probably due to the difficulty of the MCMC algorithm to explore efficiently the posterior distribution. DiffPIR predictions seem relevant but underestimated in the SISR problem.
Lastly, VBLE posterior sampling ability is shown in more detail in Fig. 7. In the left part of the figure, posterior samples are sharp and realistic, showing the interest of formulating the posterior distribution in the latent space, using a uniform distribution. The MMSE-x estimate is smoother than the samples and, hence, tends to be more faithful to the target image. Additionally, the right part of Fig. 7 provides an insight into the obtained latent diversity, using posterior samples projected on two orthogonal directions in the latent space with a high variance, that is lying in the set $\{a\odot u|u\in\{-0.5,0.5\}^{C\times M\times N}\}$ with $a$ the optimized variational parameter from eq. 11. Following the first direction, the catchlight, that is the white light reflect in the eye, goes from right to left. Following the second direction, its size increases. This illustrates the uncertainty that can be learnt within this latent uncertainty model.
Therefore, in addition to state-of-the-art point estimation results, VBLE approximate posterior distribution is relevant, as it provides accurate confidence intervals and realistic posterior samples. This validates the choice of the parametric family chosen to estimate the posterior in the latent space of CAEs. In particular, an unimodal latent distribution with a similar shape as the encoder posterior $q_{\phi}(z|x)$ shape provides a remarkably simple way to model the posterior for an image restoration task.
4.5 Ablation Study
FFHQ - SISR $\times 4$ BSD - Gaussian deblur
VBLE MMSE-x 31.18 0.2290 0.8669 24.66 0.3956 0.6433
VBLE MMSE-z 31.17 0.2468 0.8662 24.73 0.4278 0.6519
VBLEz,h MMSE-x 31.18 0.2281 0.8671 24.68 0.3952 0.6444
Table 5: Quantitative results for different variants of our method. MMSE-x and MMSE-z denote the two possible estimators defined in Sec. 3.2. For Gaussian deblur, the noise level is $\sigma=7.65/255$ and kernel deviation is $\sigma_{k}=3$.
For the experiments presented in the paper, some choices have been made. In particular, the MMSE-x estimate has been preferred to the MMSE-z estimate (see Sec. 3.2), while only $z$ has been optimized as we approximated the second latent variable $h=E_{h,\phi}(\bar{z})$ (see Sec. 3.1). In this section, in order to justify these choices, we perform quantitative and qualitative results with the MMSE-z estimate and with an optimization on $(z,h)$ (denoted by VBLEz,h). All these compared variants of the proposed method share the same restoration parameters (i.e. same $\lambda$ and $\alpha$) and use the same CAE.
Figure 8: Visual results for different variants of our method. Top: SISR $\times 4$. Bottom: Gaussian deblurring ($\sigma_{k}=3$, $\sigma=7.65/255$).
Table 5 shows quantitative results for VBLE and its variant on two inverse problems. VBLEz,h and the reference version yield almost identical results. Thus, the optimization on $(z,h)$ does not perform better than the optimization on $z$ only. For that reason, we choose to optimize only on $z$ for the main experiments, as the formulation is much simpler. Furthermore, the MMSE-z estimate yields similar PSNR and SSIM results as the MMSE-x estimate but exhibits always a poorer LPIPS. Furthermore, visual results are given in Fig. 8. While VBLEz,h and the reference version are visually very similar, the MMSE-z estimate is less textured and detailed compared to the MMSE-x, especially on the deblurring problem. This, combined with slightly lower metrics, justifies the use of the MMSE-x estimate for the main experiments.
Additionally, a study on VBLE hyperparameters $\alpha$ (bitrate parameter) and $\lambda$ (regularization parameter) is provided in Fig. 9. Parameter $\lambda$ seems a more sensitive parameter than $\alpha$, hence a CAE with a single bitrate could solve a wide range of inverse problems. Furthermore, the prior on FFHQ seems to be more informative compared to the one on BSD, which makes sense as BSD dataset structure is much harder to learn.
@@ 5 Conclusion
In this paper, we have proposed the use of variational compressive autoencoders for regularizing imaging inverse problems. Additionally, we have introduced the VBLE algorithm which performs variational inference in the latent space of CAEs or VAEs. VBLE enables efficient approximation of the posterior distribution associated with image restoration tasks. In our experiments, compressive autoencoders, combined with VBLE, achieve state-of-the-art results and provide compelling posterior sampling abilities, outperforming other posterior sampling methods in terms of computational time and GPU load. These demonstrate the interest of CAEs for image restoration as well as the possibility of estimating simply and efficiently the posterior distribution in the latent space of a generative model.